# Scalable Multi-Agent Path Finding using Collision-Aware Dynamic Alert Mask and a Hybrid Execution Strategy

### v1

@@ 5 Results and Discussion
Here we present an empirical study, evaluating our central hypothesis: a MAPF framework with strategically reduced information sharing can achieve robust performance. We address our RQs concerning minimal information needs and comparative performance against approaches with different information paradigms using the experimental results presented in Table 1.
@@
(RQ1): Given the observation constraints of a decentralized setup, can an effective MAPF algorithm be created with one agent knowing nothing about other agents? If not, what information must it need at a minimum? Ans. No—if agents have zero knowledge of one another, MAPF fails, losing completeness guarantees; but solutions are possible with targeted, minimal information sharing. Table 5 in the appendix presents comparative results for various inter-agent information sharing settings.
Now, we turn our attention to the strategically limited minimal information sharing case. In our framework, strategically informed alerts convey either (S4.1) static constraints (the conflicting cell) or (S4.2) slightly richer, yet still localized, dynamic constraints (the sub-path of only the directly conflicting agent(s) for a short horizon). Table 1 shows performance comparison of considered methodologies across 11x11 maze (a), 21x21 maze (b), and $25\times 25$ (c) warehouse maps, respectively. As shown in Table 1, the performance of our framework, where agents initially plan with only local information and receive minimal, targeted alerts from a central coordinator upon conflict, achieves high Success Rates (SR). Specifically, on the $11\times 11$ and $21\times 21$ maze grids, SRs are high (90%-100%). On the more structured $25\times 25$ warehouse map (Table 1(c)), Alert-BFS and Alert-A* also achieve 100% SR for 32 and 64 agents, with Alert-A* maintaining a 60% SR in the dense agent scenario (96 agents), demonstrating effectiveness across different environment types with this minimal information sharing strategy. This success with minimal inter-agent information contrasts sharply with alternatives. Centralized planners (CBS, ICBS) demand complete global knowledge of all agents’ paths, a requirement that, while enabling optimality, leads to their low SR due to the combinatorial explosion in conflict resolution within practical time limits. In contrast, distributed learning frameworks (PRIMAL & SCRIMP) often operate under the assumption that each planning agent possesses full observation, of other agents and their goals, in a local field of view. While this localized observation can facilitate decentralized decision-making, it implies a continuous and potentially substantial information gathering (e.g., via onboard cameras or LiDAR sensors) and processing requirement for each agent.
The success of our method across maze and warehouse environments underscores that this on-demand information is sufficient for high SR, thereby avoiding the continuous, dense local information exchange. Our quantitative analysis, detailed in the Appendix A.1, shows that this on-demand alert mechanism reduces the total information load by $\sim$93% compared to a continuous-observation distributed paradigm in this scenario.
Table 1: Comparison of performance metrics for different methods on an $11\times 11$ grid (5–20 agents, $\sim$45% obstacles), a $21\times 21$ grid (32, 64, and 96 agents, $\sim$35% obstacles), and a $25\times 25$ warehouse grid (32, 64, 96 agents, $\sim$24% obstacles). Results are averaged over 10 problem instances per configuration. Key metrics include Success Rate (SR), Makespan (MS), Collisions (CO), and Time (T), with Time averaged over all trials. For learning-based methods, the average steps taken while solving the problem instances are presented in brackets along with average wall-clock time taken.
(a) 11x11 maze with 50 sec. time limit per problem instance ($256$ steps for learning-based methods)
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (s) ($\downarrow$)
Methods 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20 5-10 11-15 16-20
Alert-BFS 100.00% 100.00% 98.00% 17.8 23.6 27.4 9 53 191 1.1 4.2 11.8
Alert-A* 100.00% 96.00% 82.00% 16.1 20.7 24.0 10 59 173 3.3 12.0 25.8
CBS 85.00% 34.00% 2.00% 13.8 14.3 17.0 559 8785 3895 8.3 30.7 35.3
ICBS 85.00% 26.00% 10.00% 14.0 14.8 12.8 885 3486 1466 5.9 25.1 28.6
PRIMAL 61.67% 48.00% 28.00% 71.7 89.9 104.8 - - - 1.4 (142) 3.4 (176) 6.1 (214)
SCRIMP 95.00% 70.00% 60.00% 19.3 28.6 38.8 1 2 8 0.4 (31) 1.6 (97) 3.9 (124)
(b) 21x21 maze with 30 min. time limit per problem instance ($512$ steps for learning-based methods)
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Alert-BFS 100.00% 90.00% 0.00% 86.6 131.4 - 528 18909 - 2.7 23.9 30.0
Alert-A* 100.00% 80.00% 0.00% 41.4 72.0 - 457 11379 - 3.3 23.5 30.0
CBS 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
ICBS 0.00% 0.00% 0.00% - - - - - - 30.0 30.0 30.0
PRIMAL 0.00% 0.00% 0.00% - - - - - - 0.8 (512) 2.1 (512) 1.8 (512)
SCRIMP 100.00% 20.00% 10.00% 65.9 135.0 94.0 15 220 357 0.1 (66) 1.0 (437) 2.4 (470)
(c) 25x25 warehouse with 1 hour time limit per problem instance ($512$ steps for learning-based methods)
SR ($\uparrow$) MS ($\downarrow$) CO ($\downarrow$) T (min) ($\downarrow$)
Methods 32 64 96 32 64 96 32 64 96 32 64 96
Alert-BFS 100.00% 100.00% 10.00% 94.3 125.9 96.0 237 6379 26553 5.2 28.0 59.5
Alert-A* 100.00% 100.00% 60.00% 40.5 43.7 50.8 211 3129 16277 4.7 19.2 51.8
CBS 0.00% 0.00% 0.00% - - - - - - 60.0 60.0 60.0
ICBS 10.00% 0.00% 0.00% 38.0 - - 123107 - - 57.9 60.0 60.0
PRIMAL 0.00% 0.00% 0.00% - - - - - - 1.0 (512) 3.1 (512) 6.6 (512)
SCRIMP 100.00% 100.00% 90.00% 60.2 49.0 101.7 2 14 161 0.1 (60) 0.3 (49) 1.1 (143)
@@
(RQ2): How does the proposed hybrid method compare to leading alternative search- and learning-based approaches in terms of performance, solution quality, and scalability? Ans. Our hybrid framework exhibits a compelling balance of performance and scalability with distinct trade-offs against established search-based and learning-based methods, as detailed in Table 1.
Our Python-based framework demonstrates higher SR and scalability compared to search-based optimal solvers. CBS/ICBS (C++ based) shows constantly lower SR with increasing environment complexity, failing to solve any instances on the $21\times 21$ maze (for 32+ agents) and the $25\times 25$ warehouse map. While CBS/ICBS achieves optimal MS when they solve (Table 1(a)), their reported Collision (CO) counts (Table 1(a)) for such cases are very high; our methods report significantly fewer collisions when they succeed on these smaller maps.
PRIMAL exhibits low SR across all tested configurations, often hitting step limits without success. SCRIMP is a strong learning-based performer with fast inference times and good MS. On the $11\times 11$ and $21\times 21$ mazes, its SR is comparable or better than our methods in some low-density cases, but can decline in denser scenarios ($21\times 21$, 64 agents: SCRIMP 20% SR vs. Alert-BFS 90% SR). On the $25\times 25$ warehouse map, SCRIMP maintains 100% SR for 32 and 64 agents, and achieves 90% SR for 96 agents, outperforming Alert-BFS (10% SR) and Alert-A* (60% SR) at this highest agent count in terms of success. However, SCRIMP also demonstrates significantly lower CO values than our methods across all successful warehouse instances attributing to its efficient multi-agent training. A critical factor for online planners is the offline training investment. For instance, PRIMAL’s extensive training on varied map sizes involved approximately 20 days on significant computational resources (Damani et al. 2021a). SCRIMP, with its sophisticated transformer-based architecture and PPO training, also implies a substantial training regimen across diverse configurations to achieve its reported performance (Wang et al. 2023). In contrast, our RL planning model, trained only on $11\times 11$ maze grids, was developed with a comparatively modest regimen: training was conducted on a Mac Mini equipped with an Apple M4 Pro chip (14-core CPU, 20-core GPU, 16-core Neural Engine) and 48GB unified memory, running for approximately 6 to 8 hours. Its subsequent effective performance when tested on different map types (map and warehouse) and larger grid sizes (up to $25\times 25$) indicates strong generalization from this simpler training setup.
@@ 6 Conclusions
Our hybrid framework combines decentralized RL-based planning with only minimal, targeted coordinated alerts–static conflict cell flags or brief conflict tracks–to maintain better SR across various challenging scenarios. While it incurs more collisions (CO) and sometimes longer makespan (MS) compared to highly specialized methods like SCRIMP, and its SR declines in extremely dense scenarios, it often delivers better overall feasibility and scalability, in scenarios with increasing agent counts, than purely search‑based methods and PRIMAL by strategically sharing minimum inter-agent information.
Effective multi-agent coordination with reduced information exchange, as explored in this research, will lead to privacy-aware automated systems, enabling broader use in complex privacy-preserving applications like autonomous driving. Despite these promising results, our work has limitations. The current framework, optimized for disappear at target setting, may face challenges in stay at target behavior. Future work will investigate the minimal information necessary to effectively manage these stay at target scenarios, aiming to extend the framework’s applicability without compromising its core principle of sparse information exchange. Additionally, an ablation study of richer AlertMask encodings, like dynamic occupancy maps, could improve replanning efficiency.
@@ Appendix A Appendix
A.1. A Quantitative Model of Information Load ........................................................................................................................................................................A.1
A.2. Literature review ........................................................................................................................................................................A.2
    Literature Categorization Table 2 ........................................................................................................................................................................2
A.3. Training Methods ........................................................................................................................................................................A.3
  A.3. PPO Training Procedure ........................................................................................................................................................................A.3
  A.3. Neural Network Architecture ........................................................................................................................................................................A.3
  A.3. Model Comparison ........................................................................................................................................................................A.3
@@
  A.3. Model Comparison ........................................................................................................................................................................A.3
    Table 3 ........................................................................................................................................................................
A.4. Additional Experiments ........................................................................................................................................................................A.4
  A.4. Agent Selection Policy Comparison ........................................................................................................................................................................A.4
    Table 4 ........................................................................................................................................................................4
  A.4. Inter-agent Information Sharing Comparison ........................................................................................................................................................................A.4
    Table 5 ........................................................................................................................................................................5
  A.4. Benchmark Dataset Evaluation ........................................................................................................................................................................A.4
    Table 6 ........................................................................................................................................................................6
A.1 A Quantitative Model of Information Load
To quantitatively support our claim of reduced information sharing, we developed a simplified model to estimate the ”Total Information Load” of different coordination paradigms. This model is based on a simple Information Unit (IU), defined as the data required to represent one agent’s state (e.g., its coordinates) at a single timestep.
Model Assumptions
Our calculation is based on the 11x11 maze scenario with 20 agents, using metrics from the successful Alert-BFS runs in Table 1(a).
  1. 1.
    Scenario Data: We use a 20-agent instance (N=20) solved with a makespan (T) of 27 timesteps and requiring the resolution of 191 conflicts (C).
  2. 2.
    Average Path Length (L): We assume an average initial path length of 25 steps per agent.
  3. 3.
    Agent Density (D): For a distributed method like PRIMAL with a 10x10 FoV on an 11x11 grid, each agent has near-global vision. Therefore, each agent continuously observes all other N-1 agents. We set the average density (D) to 19 neighbors.
  4. 4.
    Alert Size ($I_{\text{alert}}$): We assume each conflict alert sent by our coordinator is a minimal, static alert costing 1 IU.
Comparative Calculation
Distributed Method (e.g., PRIMAL)
The information load is the continuous sensing and processing of all visible neighbors by every agent at every timestep. The formula is $\text{Info}_{\text{Distributed}}=N\times T\times D$.
$\text{Info}_{\text{Dist.}}$ $=20\times 27\times 19$
$=\textbf{10,260 IU}$
Our Hybrid Method (Alert-X)
The load is the one-time path sharing in addition to all subsequent targeted alerts. The formula is $\text{Info}_{\text{Alert-X}}=(N\times L)+(C\times I_{\text{alert}})$.
$\text{Info}_{\text{Alert-X}}$ $=(20\times 25)+(191\times 1)$
$=500+191$
$=\textbf{691 IU}$

### v2

@@ 5 Results and Discussion
Here we present an empirical study, evaluating our central hypothesis: a MAPF framework with strategically reduced information sharing can achieve robust performance. We address our RQs concerning minimal information needs, and Table 3 summarizes the comparative performance of IC–MAPF against representative search-based (CBS) and learning-based (DCC, SCRIMP, EPH) MAPF solvers across random grids and structured maps. Additional experimental results are presented in Appendix Section A.1 and detailed results with runtime in Figure 3.
@@
(RQ1): Given the observation constraints of a decentralized setup, can an effective MAPF algorithm be created with one agent knowing nothing about other agents? If not, what information must it need at a minimum?
Ans. No–A purely decentralized setting in which agents have no information about others is insufficient once interactions occur, as conflicts cannot be consistently resolved. However, our results show that effective MAPF does not require continuous observation or global plan sharing. Instead, solutions can be achieved with minimal, event-triggered information sharing consistent with our hybrid information model $\mathcal{I}_{i}^{\text{hybrid}}(t)$, where additional other-agent information is revealed only upon conflict events.
Across unstructured random grids, IC–MAPF achieves high success rates while using orders of magnitude less information than learning-based baselines. On the random-$\times32$-20 map, IC–MAPF maintains $100\%$ SR up to 80 agents, with cumulative IU values ranging from 3.6 to 214.2. Under the same settings, SCRIMP requires 262–2823 IU and EPH requires 19.7–6155 IU, corresponding to approximately $3\times$ to over $30\times$ higher information usage, while DCC requires 26.1–6337.2 IU (about $7\times$$30\times$ IC–MAPF over the same range). On the larger random-$\times64$-20 map, IC–MAPF sustains $100\%$ SR up to 80 agents with IU below 212, whereas SCRIMP fails beyond 30 agents and exceeds 6000 IU, and EPH exhibits reduced SR while using $6\times$ to $13\times$ more information than IC–MAPF; DCC similarly achieves moderate SR but uses 10.6–2947.4 IU, i.e., roughly $7\times$$14\times$ more information. Similarly, on structured benchmarks, den312d, IC–MAPF achieves $100\%$ SR for 8–64 agents with IU between 3.6 and 759.9, while EPH requires up to 6722.3 IU to obtain comparable SR, and SCRIMP fails for 16 agents and above; DCC lies in between, using 22.6–8180.1 IU (approximately $6\times$$11\times$ IC–MAPF for 8–64 agents). On the warehouse map, IC–MAPF solves all tested instances (8–128 agents) with IU between 0.8 and 139.1, whereas EPH requires 1.7–1502.4 IU to achieve similar SR, and DCC uses 4.3–2016.5 IU (about $5\times$$16\times$ IC–MAPF), while SCRIMP achieves $0\%$ SR despite using several thousand IU. These results identify targeted, conflict-triggered sharing of short agent-path fragments over bounded horizons as the minimal information required for effective MAPF, enabling reduced total information load of up to one to two orders of magnitude.
@@
(RQ2): How does the proposed hybrid method compare to leading alternative search- and learning-based approaches in terms of performance, solution quality, and scalability?
Ans. Table 3 comparison highlights three key dimensions: success rate (SR), makespan (MS), and total information load (IU). Search-based solvers such as CBS rely on full joint information $\mathcal{I}_{i}^{\text{central}}(t)$ and attempt to resolve all couplings through centralized reasoning. While optimal when successful, CBS fails to scale under the imposed time limits and does not solve most medium-to-large instances reported in the table. This illustrates the practical limitations of centralized planners as problem size grows. Learning-based distributed methods, including DCC, SCRIMP, and EPH, operate under $\mathcal{I}_{i}^{\text{distributed}}(t)$, where each agent continuously accesses local observations of other agents within a field of view or via explicit message passing. These methods show strong performance in some dense scenarios (e.g., EPH on warehouse), but incur substantially higher information loads. In Table 3, their IU values are reported as multipliers relative to IC–MAPF, often ranging from one to two orders of magnitude higher due to continuous perception or communication.
In contrast, IC–MAPF adopts a hybrid execution strategy. This allows IC–MAPF to scale favorably across both random and structured maps, achieving high SR with competitive makespans while keeping information exchange sparse and bounded. The gradual escalation of information, from static constraints to short dynamic sub-paths, and finally to bounded joint replanning only when necessary, ensures that increased coordination cost is incurred only in genuinely hard cases. Overall, IC–MAPF occupies a middle ground between centralized optimal planners and fully distributed learning-based planners. These results demonstrate that carefully controlled, event-driven information sharing can provide a favorable trade-off between performance and communication complexity in MAPF.
@@ 6 Conclusions
In this paper, we introduced the problem of Information-centric MAPF and proposed a hybrid framework, IC–MAPF, that significantly reduces the inter-agent information required for successful multi-agent coordination. We further defined a metric, Information Units (IU), to quantify the amount of other-agent information consumed during planning. Across a range of challenging MAPF benchmarks, IC–MAPF achieves high success rates while reducing information usage by $2\times$ to $23\times$ compared to state-of-the-art communication-efficient baselines.
The broader implications of this work extend to privacy-aware autonomous systems, where reducing inter-agent information exchange is essential. By showing that multi-agent coordination is achievable with sparse, event-triggered information, our results support the development of scalable, bandwidth-efficient, and privacy-preserving multi-robot systems. Despite these promising results, our work has limitations: IC–MAPF may face challenges in extremely dense settings where conflicts are frequent, and its tiered strategy is currently hand-designed. As future work, we aim to explore learning-based mechanisms that automatically adapt the tier-selection policy and investigate reinforcement learning under explicit information constraints, enabling agents to optimize performance while minimizing IU usage.
@@ Appendix A Appendix
A.1. Additional Results ........................................................................................................................................................................A.1
  Detailed Performance Table – Table 3 ........................................................................................................................................................................3
  Strategy Usage Summary – Figure 2 ........................................................................................................................................................................2
  Detailed Performance Plots – Figure 3 ........................................................................................................................................................................3
A.2. Properties of Tiered Replanning Strategies ........................................................................................................................................................................A.2
A.3. Hardware Results: TurtleBot4 ........................................................................................................................................................................A.3
  Robot Grid Configurations – Figure 4 ........................................................................................................................................................................4
  Physical Setup Image – Figure 5 ........................................................................................................................................................................5
  Communication Statistics – Table 4 ........................................................................................................................................................................4
  Execution Quality Statistics – Table 5 ........................................................................................................................................................................5
A.4. Literature review ........................................................................................................................................................................A.4
    Literature Categorization Table 6 ........................................................................................................................................................................6
A.5. Training Methods ........................................................................................................................................................................A.5
  A.5. PPO Training Procedure ........................................................................................................................................................................A.5
  A.5. Neural Network Architecture ........................................................................................................................................................................A.5
  A.5. Model Comparison ........................................................................................................................................................................A.5
@@
A.1 Additional Results
Tables 3, Figures 2, and Figure 3 provide expanded performance analyses complementing the compact results presented in the main paper. Table 3 reports the full per-configuration statistics for all solvers across four benchmark maps, including success rate, makespan, and absolute values for Information Units (IU). These detailed values highlight the consistency of IC–MAPF across increasing agent densities and show the scale at which learning-based methods begin to degrade or saturate, especially in larger and more structured environments.
Figure 2 summarizes the usage frequency of our tiered replanning strategies S4.0 to S4.3 (Yield, Static, Dynamic, and Joint) across all problem instances. This visualization illustrates how IC–MAPF automatically escalates strategy complexity only when needed, typically in high-density or tightly coupled regions, while relying predominantly on low-information repairs in easier configurations.
Figure 3 presents aggregate trends for success rate, makespan, IU consumption, and average runtime. The shaded regions indicate standard deviation across 20 problem instances, providing insight into solver stability and variance across different metrics. Compared to CBS and learning-based planners (DCC, SCRIMP, EPH), IC–MAPF achieves consistently high success rates while maintaining substantially lower information usage and competitive makespan behavior. Together, these additional results reinforce the main claims of the paper: IC–MAPF offers a favorable balance between scalability, robustness, and dramatically reduced inter-agent information requirements.
Search-based Hybrid Learning-based Solvers
Map $m$ CBS IC–MAPF SCRIMP DCC EPH
[…]
[Uncaptioned image] random $\times32$-20 10 100 46.8 100 50.5 3.6 100 48.0 262.2 100 49.1 26.1 100 47.6 19.7
20 100 47.6 100 56.0 18.8 100 50.5 565.2 100 58.0 211.2 95 62.5 204.2
30 100 48.4 100 55.5 66.3 100 52.4 884.8 100 59.5 696.2 100 62.0 576.8
40 90 80.0 100 107.0 48.8 95 58.0 1377.7 95 72.8 1407.0 100 64.5 1403.0
50 65 85.3 100 61.4 221.4 100 56.8 1645.2 90 81.7 3356.1 95 79.0 3183.2
60 40 90.6 100 128.2 214.2 90 71.4 2822.8 70 92.0 6337.2 95 88.7 6155.1
70 0 - 100 65.5 2158.2 90 70.5 3158.8 60 104.5 10385.0 65 110.7 11735.3
80 0 - 100 71.8 1405.5 80 83.8 4660.8 60 111.7 17196.1 70 108.9 20330.8
90 0 - 95 82.8 1909.0 60 96.9 6292.6 30 117.8 29474.4 25 123.8 36040.6
100 0 - 85 76.5 3553.8 60 97.2 6982.9 40 117.5 37877.7 50 121.0 41383.7
128 0 - 0 - - 70 96.6 8615.5 0 128.0 - 5 127.5 129548.7
[Uncaptioned image] random $\times64$-20 10 100 98.0 100 108.0 0.9 50 112.5 612.1 100 97.7 10.6 95 97.3 7.8
20 90 108.8 100 115.6 6.2 10 126.5 1382.0 100 107.8 59.4 100 105.7 47.8
30 65 112.0 100 118.2 23.4 10 126.7 2190.3 90 112.2 172.2 100 109.5 141.8
40 60 128.5 100 122.6 45.0 0 128.0 - 100 112.8 382.4 100 112.3 328.9
50 40 127.0 100 124.0 109.8 0 128.0 - 85 116.8 889.7 100 113.8 638.0
60 25 134.5 100 130.2 110.8 0 128.0 - 65 120.8 1258.0 95 116.1 964.6
70 0 - 100 128.2 183.0 0 128.0 - 50 123.5 1823.1 85 120.8 1640.2
80 0 - 100 128.7 211.9 0 128.0 - 20 126.8 2947.4 90 120.2 2520.7
90 0 - 95 130.4 245.4 0 128.0 - 60 124.6 4095.8 80 122.2 3187.4
100 0 - 90 130.1 462.1 0 128.0 - 25 127.2 6134.2 45 125.3 5290.0
128 0 - 5 132.0 607.6 0 128.0 - 0 128.0 - 15 127.7 12250.5
[…]
16 85 128.7 100 124.0 21.8 0 256.0 - 100 130.7 191.3 100 122.2 128.4
32 20 148.0 100 136.7 105.2 0 256.0 - 85 163.5 1195.6 100 127.0 627.5
64 0 - 100 153.1 759.9 0 256.0 - 90 190.7 8180.1 100 146.1 6722.3
128 0 - 0 - - 0 256.0 - 5 254.8 112519.4 90 207.7 77108.9
[Uncaptioned image] warehouse $1×\times6$3 8 100 175.7 100 192.8 0.8 0 256.0 - 100 191.8 4.3 100 191.6 1.7
16 100 196.2 100 215.1 1.0 0 256.0 - 100 213.3 15.6 100 212.9 11.1
32 15 221.3 100 220.3 10.7 0 256.0 - 100 219.2 89.0 100 218.1 57.7
64 0 - 100 227.6 23.6 0 256.0 - 100 227.6 356.9 100 227.8 265.1
128 0 - 100 234.7 139.1 0 256.0 - 95 236.8 2016.5 90 239.1 1502.4
[…]
Figure 2: Strategy usage summary on random-32-32-20 (random-32), random-64-64-20 (random-64), den312d, and warehouse maps. Values are averaged across 20 problem instances.
[…]